solution containing the complex formed between bi( iii) and thiourea has a molar absorptivity of 9.32 x 103 l mol-1 cm at 470 nm_ (a) what is the absorbance of a 6.24xl0 5 m solution of the complex at 470 nm in a 1.00 cm cell? (b) what is the percent transmittance of the solution described in (a)? (c) what is the molar concentration of the complex in solution that has the absorbance described in

1
Expert's answer
2022-04-05T14:03:03-0400

C

=

2

×

10

4

M


A

=

0.123


l

=

2

c

m


?

T

, ?

%

T

, ?

ε

The relationship between transmittance T

 and absorbance is:

A

=

log

T

T

=

10

A

Since the length of the cell was increased to l

=

2

c

m

, the absorbance will double and therefore, A

2

=

2

×

0.123

=

0.246

T

=

10

0.246

=

0.568

Therefore, %

T

=

100

×

T

=

100

×

0.568

=

56.8

%

In order to solve for the molar absorptivity ε

, we will use the Beer-Lambert law:

A

=

ε

×

l

×

C

ε

=

A


l

×

C

ε

=

0.123


1

c

m

×

2

×

10

4

M

=

615

M

1

c

m

1


The basic idea here is to use a graph plotting Absorbance vs. Concentration of known solutions. Once you have that you can compare the absorbance value of an unknown sample to figure out its concentration.

http://intro.chem.okstate.edu/ChemSource/Instrument/inst4.htm

You will be applying Beer's law to calculate the concentration.

The equation for Beer's law is: A = εmCl

(A=absorbance, εm = molar extinction coefficient, C = concentration, l=path length of 1 cm)

You should have a data set which was used to create a standard curve. The graph should plot concentration (independent variable) on the x-axis and absorption (dependent variable) on the y axis.

You'll need to add a line of best fit to the data points and determine the equation for the line. The equation should be in y=mx + b form.

y = absorbance (A)

Note: no unit for absorbance

x = concentration (C)

Note: unit is M or mol/L

m = (εm) = slope or the molar extinction coefficient in beers law which has units of M

1

c

m

1

So A = εmC +b

If you solve for C you should get

C = (A-b)/εm